OpenStudy (anonymous):

I have to find the slope of 3x+4y=22. I think I have part of the problem.. but i am stuck

OpenStudy (anonymous):

3x+4y=22 -3x=22-3x 4y=22-3x is this correct so far?

OpenStudy (anonymous):

you have to isolate for y in order to get the slope

OpenStudy (anonymous):

so you would get y= 3/4x+22/4 and the slope is the coefficient of a which is 3/4

OpenStudy (anonymous):

Thanks...

OpenStudy (anonymous):

Thanks Lemonhead71... Do you just come in to save the day?